OpenStudy (gabbyalicorn):

Can someone double check this? Thanks! A hard candy has a radius of 9 mm. The candy has a spherical soft center with a radius of 6 mm, with a hard shell surrounding it. What is the volume of the hard shell? Use 3.14 to approximate pi and express your final answer in hundredths. I have 1981.39 is this correct?

OpenStudy (welshfella):

you used the formula for the volume of a sphere right? V = (4/3) pi r^3

OpenStudy (gabbyalicorn):

Yes.

OpenStudy (gabbyalicorn):

V = (4/3)π(r)^3 V = (4/3)3.14(9-6)^3 V = (4/3)3.14(3)^3 V = 4/3(9.42)^3 V = 12.56^3 V = 1981.385216

OpenStudy (welshfella):

Ok so V hard shell = (4/3) *3.14( 9^3 - 6^3)

OpenStudy (gabbyalicorn):

Which would equal 2147.76

OpenStudy (welshfella):

right

OpenStudy (welshfella):

(9 - 6)^3 was not correct

OpenStudy (gabbyalicorn):

ok

OpenStudy (welshfella):

yw

OpenStudy (gabbyalicorn):

Tysm.